The equilibrium constant, Kc, is calculated using molar concentrations. For gaseous reactions another form of the equilibrium constant, Kp, is calculated from partial pressures instead of concentrations. These two equilibrium constants are related by the equation Kp=Kc(RT)Δn where R=0.08206 L⋅atm/(K⋅mol), T is the absolute temperature, and Δn is the change in the number of moles of gas (sum moles products - sum moles reactants). For example, consider the reaction N2(g)+3H2(g)⇌2NH3(g) for which Δn=2−(1+3)=−2.

A) For the reaction 3A(g)+3B(g)⇌C(g) Kc = 78.0 at a temperature of 211 ∘C . Calculate the value of Kp.

B) For the reaction X(g)+3Y(g)⇌3Z(g) Kp = 2.83×10−2 at a temperature of 171 ∘C . Calculate the value of Kc .
Answer #1

A)

Kc = 78.0

Now, n = 1 - (3+3) = -5

T = 211 +273= 484 K

So, Kp = Kc × (RT)n

Or, Kp = 78.0 × (0.08206× 484)-5

Or, Kp = 7.89 × 10-7 .

B)

Kp = Kc × (RT)n

Or, Kc = Kp × (RT)- n

For the given reaction

n = 3 - (1+3) = -1.

T = 171+273= 444 K

So,

Kc = Kp × ( 0.08206 × 444)1

Or, Kc = 2.38 × 10-2 × (0.08206×444)

Or, Kc = 0.867 .